### v1

# Resolvent estimates for a function of a linear operator

Abstract.
Let $T$ be a bounded linear operator on a Banach space and $f$ an analytic function, defined on the spectrum of $T$. We discuss the relations between the rate of growth of the resolvent of $T$ and of $f(T)$.
Key words and phrases:
Resolvent growth
2020 Mathematics Subject Classification:
47A10, 47A30
Both authors acknowledge the support of the Spanish Ministry for Science and Innovation under Grant PID2022-137294NB-I00, DGI-FEDER. G. Bello has also been partially supported by PID2022-138342NB-I00 for Functional Analysis Techniques in Approximation Theory and Applications (TAFPAA) and by Project E48_23R, D.G. Aragón
1. Introduction
Let $T$ be a bounded linear operator on a Banach space $X$. As usual, let $\sigma(T)$ denote the spectrum of $T$. Given a compact set $K$ containing $\sigma(T)$ and a real number $s\geq 1$, we say that $T$ has resolvent growth of order $s$ near $K$ if
$\lVert(T-\lambda)^{-1}\rVert\leq C\operatorname{dist}(\lambda,K)^{-s},\quad\lambda\in D(a,R)\setminus K,$ (1.1)
where $C$ is a constant and $D(a,R)=\{\lambda\in\mathbb{C}:|\lambda-a|<R\}$ is an open disc containing $K$. The growth behavior is really determined by those $\lambda$ close to $K$, so this definition does not depend on the choice of $D(a,R)$. An equivalent way to write (1.1) is
$\lVert(T-\lambda)^{-1}\rVert\leq C^{\prime}\max\{\operatorname{dist}(\lambda,K)^{-s},\operatorname{dist}(\lambda,K)^{-1}\},\quad\lambda\in\mathbb{C}\setminus K.$
When $K=\sigma(T)$ it is said that $T$ has resolvent growth of order $s$.
Let $\Omega$ be an open set containing $K$ and let $f\colon\Omega\to\mathbb{C}$ an analytic function. This paper is devoted to the study of the relation between the resolvent growth of $T$ near $K$ and the resolvent growth of $f(T)$ near $f(K)$, the latter operator being defined by the Riesz–Dunford calculus. Consequently, we will assume throughout this paper that $f$ is not constant on any connected component of $\Omega$. Notice also that $\Omega$ can be assumed as small as needed (containing $K$), so in particular we can suppose that $\Omega$ is bounded.
The most relevant case in our results is when the compact set $K$ is precisely $\sigma(T)$. We work with a general $K$ because the proofs are just the same.
Theorem A (Bakaev [2]).
If $T$ has resolvent growth of order $s$ near $K$, then $f(T)$ also has resolvent growth of order $s$ near $f(K)$.
Bakaev formulated this theorem for rational functions $f$. We will give a complete proof for analytic functions.
By a Lipschitz curve in $\mathbb{C}$ we mean a curve that can be parametrized as $z=\zeta\cdot(t+if(t))$, $t\in I$, where $I$ is a finite interval, $f\colon I\to\mathbb{R}$ a Lipschitz function and $\zeta$ a complex constant of modulus $1$. Here we will prove the following partial converse of this result.
For a set $G\subset\mathbb{R}^{2}$ and $\sigma>0$, we put
$G]_{\sigma}:=\{y\in\mathbb{R}^{2}\setminus G:\;0<\operatorname{dist}(y,G)\leq\sigma\}.\quad$
Let $p\in(0,2)$ be a real number and let $G\subset\mathbb{C}$ be a compact set. We say that $G$ is $p$-admissible if there is a constant $C$ such that for any $a\in\mathbb{R}^{2}$ and any $R,\sigma>0$,
$m\big{(}[G]_{\sigma}\cap D(a,R)\big{)}\leq C\sigma^{2-p}R^{p}.$ (1.2)
It is clear that each Lipschitz curve is $1$-admissible, but not vice versa. If the Assouad dimension of a compact set is less than two, then this set is $p$-admissible for some $p<2$. We refer to [3] for details.
Theorem 1.1.
Let $T$ be a Banach space operator. Suppose $f(T)$ has resolvent growth of order $s$ near $f(K)$, where $K\supset\sigma(T)$, and that $f^{\prime}$ does not vanish on $K$.
  • (i)
    If $K$ is contained in a finite union of Lipschitz curves, then $T$ has resolvent growth of order $s$ near $K$.
  • (ii)
[…]
The statement of Theorem 1.1 is false if $K=\sigma(T)$ and $f^{\prime}$ has zeros on $\sigma(T)$ (just take as $T$ a $2\times 2$ Jordan block, whose diagonal element is a zero of $f^{\prime}$). We will see in Example 3.2 that the statement of Theorem 1.1 may fail if $K=\sigma(T)$ and $\sigma(T)$ does not satisfy the above geometric condition.
Theorem 1.1 will rely on the following fact, which will be deduced from the results of [3].
Lemma 1.2.
Let $T$ be a Banach space operator and $K,\widetilde{K}\subset\mathbb{C}$ be compact sets such that $\sigma(T)\subset K\subset\widetilde{K}$ and
$\lVert(T-\lambda)^{-1}\rVert\leq C\operatorname{dist}(\lambda,\widetilde{K})^{-s},\quad\lambda\in D(b,R)\setminus\widetilde{K}$, (1.3)
where $D(b,R)$ is an open disc containing $\widetilde{K}$.
  • (i)
    If $\widetilde{K}$ is contained in a finite union of Lipschitz curves, then $T$ has resolvent growth of order $s$ near $K$.
  • (ii)
[…]
We recall that $T_{0}\in L(X)$ is called scalar spectral if there is a projection-valued spectral measure $E(\cdot)$, defined on all Borel subsets of $\mathbb{C}$, such that
$T_{0}=\int_{\mathbb{C}}\lambda\,dE(\lambda).$
In the Hilbert space case, this is equivalent to $T_{0}$ being similar to a normal operator. An operator $T\in L(X)$ is called spectral if it has a representation $T=T_{0}+N$, where $T_{0}$ is scalar spectral and $N$ is a quasinilpotent operator such that $NE(\delta)=E(\delta)N$ for any Borel set $\delta\subset\mathbb{C}$. We refer to [4] for the basics on spectral operators and to [6] for their simplified definition.
Each scalar spectral operator has first order growth of the resolvent. More generally, if $T$ is scalar and its quasinilpotent part $N$ satisfies $N^{k}=0$ for some $k\in\mathbb{N}$ (i.e., $N$ is nilpotent), then the resolvent of $T$ has $k$th order growth. Therefore, the following result can be seen as a counterpart of the above Theorem 1.1.
Theorem B (Apostol [1]).
If $f(T)$ is (scalar) spectral and $f^{\prime}$ does not vanish on $\sigma(T)$, then $T$ is (scalar) spectral.
The questions posed here are certainly interesting for unbounded operators. In particular, the case when $T$ is an (unbounded) generator of a $C_{0}$ group and one has to relate the resolvent growth of $T$ and that of the operators in the group $\{\exp(sT)\}$ for a single value of $s$ (or for all values $s\in\mathbb{R}$). We refer to Chapter 11 of the book [5] by S. Kantorovitz for information on this problem in some special setting.
2. Lemmas on analytic functions
Fix, as in the Introduction, a bounded linear operator $T$ on a Banach space $X$, a bounded open set $\Omega$ containing $\sigma(T)$, and an analytic function $f\colon\Omega\to\mathbb{C}$ that is not constant in any connected component of $\Omega$.
Then $f^{\prime}$ has only finitely many zeros on $K$, which we denote by $\xi_{1},\dots,\xi_{\ell}$. We will assume that $f$ is bounded on $\Omega$ and that $f^{\prime}$ has no zeros on $\Omega\setminus K$; this is achieved by substituting $\Omega$ with a smaller open set $\Omega^{\prime}$ such that $K\subset\Omega^{\prime}\subset\Omega$. The orders $m_{j}$ of the branching points $\xi_{j}$ of the mapping $f$ satisfy $m_{j}\geq 2$. By basic complex analysis, there is a small $\varepsilon>0$ and neighbourhoods $W_{j}$ of $\xi_{j}$ with disjoint closures contained in $\Omega$ such that $f$ maps $m_{j}$-to-one $W_{j}\setminus\{\xi_{j}\}$ to the punctured discs $0<|z-f(\xi_{j})|<\varepsilon$. We may also assume that the discs $D(f(\xi_{j}),\varepsilon)$, $D(f(\xi_{k}),\varepsilon)$ are disjoint whenever $f(\xi_{j})\neq f(\xi_{k})$ and that
${D}^{\text{\,\rm cl}}(\xi_{k},2\operatorname{diam}W_{k})\subset\Omega.$ (2.1)
We put $\Omega_{\text{br}}=\bigcup_{j}W_{j}$. Consider the compact set $(K\setminus\Omega_{\text{br}})\cup\partial\Omega_{\text{br}}$. Each $\lambda$ in $(K\setminus\Omega_{\text{br}})\cup\partial\Omega_{\text{br}}$ has a neighborhood $U$ in $\Omega$ such that ${U}^{\text{\,\rm cl}}\subset\Omega$, $f$ is one-to-one in $U$, $f^{\prime}\neq 0$ in ${U}^{\text{\,\rm cl}}$, the image $V=f(U)$ is open, and the function $g\colon V\to U$ defined by $f(g(z))=z$ is analytic in $V$. In this situation, we say that $g$ is an analytic branch of $f^{-1}$. We will use the notation $U\leftrightarrow V$ to express that there exists an analytic branch of $f^{-1}$ from $V$ to $U$.
For each $\lambda\in(K\setminus\Omega_{\text{br}})\cup\partial\Omega_{\text{br}}$ consider a neighborhood $U=U_{\lambda}$ as above. Without loss of generality, assume that $U_{\lambda}$ is an open disc. Let $U_{\lambda_{1}},\dotsc,U_{\lambda_{d}}$ be a finite cover of $(K\setminus\Omega_{\text{br}})\cup\partial\Omega_{\text{br}}$. Let $\Omega_{1}$ and $\Omega_{2}$ be open sets such that
$(K\setminus\Omega_{\text{br}})\cup\partial\Omega_{\text{br}}\subset\Omega_{2}\subset{\Omega}^{\text{\,\rm cl}}_{2}\subset\Omega_{1}\subset{\Omega}^{\text{\,\rm cl}}_{1}\subset U_{\lambda_{1}}\cup\dotsb\cup U_{\lambda_{d}}.$ (2.2)
We will also set
$\Omega_{1br}=\Omega_{1}\cup\Omega_{\text{br}},\quad\Omega_{2br}=\Omega_{2}\cup\Omega_{\text{br}}.$
Notice that $K\subset\Omega_{2br}\subset\Omega_{1br}$.
The following result simply says that $f$ is Lipschitz on the compact set ${\Omega}^{\text{\,\rm cl}}_{1br}$, which follows from well-known facts. We give the proof for the sake of completeness.
Lemma 2.1.
There is a constant $C$ such that
$\lvert f(\lambda)-f(\widetilde{\lambda})\rvert\leq C\lvert\lambda-\widetilde{\lambda}\rvert$
for every $\lambda,\widetilde{\lambda}\in\Omega_{1br}$.
Proof.
Let $D_{1},\dotsc,D_{n}$ be open discs such that
[…]
and let $\delta>0$ be a Lebesgue number associated to this open covering, that is, if $A\subset\Omega_{1br}$ has diameter $\operatorname{diam}A<\delta$, then $A\subset D_{j}$ for some $j$. Take $\lambda,\widetilde{\lambda}\in\Omega_{1br}$. If $\lvert\lambda-\widetilde{\lambda}\rvert\geq\delta$, then
$\lvert f(\lambda)-f(\widetilde{\lambda})\rvert\leq\operatorname{diam}f(\Omega_{1br})\leq\frac{\operatorname{diam}f(\Omega_{1br})}{\delta}\lvert\lambda-\widetilde{\lambda}\rvert=:C_{1}\lvert\lambda-\widetilde{\lambda}\rvert$.
On the other hand, if $\lvert\lambda-\widetilde{\lambda}\rvert<\delta$, then both $\lambda$ and $\widetilde{\lambda}$ belong to some disc $D_{j}$. By the convexity of discs, an application of the mean value theorem gives that $\lvert f(\lambda)-f(\widetilde{\lambda})\rvert\leq C_{2}\lvert\lambda-\widetilde{\lambda}\rvert$, where
$C_{2}:=\max\{\lvert f^{\prime}(s)\rvert:s\in{(D_{1}\cup\dotsb\cup D_{n})}^{\text{\,\rm cl}}\}.$
Therefore the statement follows by taking $C:=\max\{C_{1},C_{2}\}$. ∎
Lemma 2.2.
There is a positive number $\delta$ such that any two points $\lambda\neq\lambda^{\prime}$ in $\Omega_{1br}$ that satisfy $f(\lambda)=f(\lambda^{\prime})$ and $|\lambda-\lambda^{\prime}|<\delta$ belong to the same set $W_{j}$, for some $j$.
Proof.
Let $\delta$ be a Lebesgue number associated to the open covering
$\Omega_{1br}\subset U_{\lambda_{1}}\cup\dotsb\cup U_{\lambda_{d}}\cup W_{1}\cup\dotsb W_{\ell}.$
Then, whenever $\lambda,\lambda^{\prime}$ satisfy the above conditions, these points belong to the same set of the open covering. This set cannot be a set $U_{\lambda_{j}}$, because $f$ is injective on each of these sets. So both $\lambda$ and $\lambda^{\prime}$ belong to a set $W_{j}$ for some $j$. ∎
Notice that $f^{\prime}\neq 0$ on ${\Omega}^{\text{\,\rm cl}}_{1}$. We set
$\tau:=\min\{\lvert f^{\prime}(\lambda)\rvert:\lambda\in{\Omega}^{\text{\,\rm cl}}_{1}\}>0.$ (2.3)
Lemma 2.3.
There is a positive number $\rho$ such that every $\lambda$ in ${\Omega}^{\text{\,\rm cl}}_{2}$ has a neighborhood $Y_{\lambda}$ in $\Omega_{1}$ satisfying $Y_{\lambda}\leftrightarrow D(f(\lambda),\rho)$.
Proof.
Let $\delta_{1}>0$ be the Lebesgue number associated to the cover $\{U_{\lambda_{j}}\}$ of ${\Omega}^{\text{\,\rm cl}}_{1}$. Take $\delta_{2}\in(0,\delta_{1}/2)$ sufficiently small so that the discs $D(\lambda,\delta_{2})$ are contained in $\Omega_{1}$ for all $\lambda\in{\Omega}^{\text{\,\rm cl}}_{2}$. We claim that $\rho:=\tau\delta_{2}/4$ satisfies the statement. Indeed, fix $\lambda\in{\Omega}^{\text{\,\rm cl}}_{2}$. Since $\operatorname{diam}D(\lambda,\delta_{2})<\delta_{1}$, we have $D(\lambda,\delta_{2})\subset U_{\lambda_{j}}$ for some $j\in\{1,\dotsc,d\}$. By the Koebe $1/4$ theorem, $D(f(\lambda),\rho)\subset f(D(\lambda,\delta_{2}))$. Therefore the set
$Y_{\lambda}:=f^{-1}(D(f(\lambda),\rho))\cap D(\lambda,\delta_{2})$
fulfills the desired properties. ∎
Given a set $S$ and functions $\alpha,\beta\colon S\to[0,\infty)$, we write
$\alpha(s)\lesssim\beta(s),\quad s\in S$
if there exists a constant $C>0$, which does not depend on $s$, such that
$\alpha(s)\leq C\beta(s)\quad\textnormal{for all }s\in S.$
If we have both $\alpha(s)\lesssim\beta(s)$ and $\beta(s)\lesssim\alpha(s)$, then we write
$\alpha(s)\asymp\beta(s),\quad s\in S.$
Put
$\widetilde{K}:=f^{-1}\big{(}f(K)\big{)}\cap{\Omega}^{\text{\,\rm cl}}_{1br}.$
Clearly, $\widetilde{K}$ is compact, $\widetilde{K}\supset K$ and $f(\widetilde{K})=f(K)$.
Lemma 2.4.
We have
$\operatorname{dist}(\lambda,\widetilde{K})\asymp\operatorname{dist}(f(\lambda),f(K)),\quad\lambda\in{\Omega}^{\text{\,\rm cl}}_{2}.$
Proof.
Let us first prove that
$\operatorname{dist}(f(\lambda),f(K))\lesssim\operatorname{dist}(\lambda,\widetilde{K}),\quad\lambda\in{\Omega}^{\text{\,\rm cl}}_{2}.$ (2.4)
For each $\lambda\in{\Omega}^{\text{\,\rm cl}}_{2}$, let $\widetilde{\lambda}\in\widetilde{K}$ be such that $\operatorname{dist}(\lambda,\widetilde{K})=\lvert\lambda-\widetilde{\lambda}\rvert$. By Lemma 2.1,
$\lvert f(\lambda)-f(\widetilde{\lambda})\rvert\lesssim\lvert\lambda-\widetilde{\lambda}\rvert,\quad\lambda\in{\Omega}^{\text{\,\rm cl}}_{2}.$
Since the left hand side above is at least $\operatorname{dist}(f(\lambda),f(K))$, (2.4) follows.
Now let us see that
$\operatorname{dist}(\lambda,\widetilde{K})\lesssim\operatorname{dist}(f(\lambda),f(K)),\quad\lambda\in{\Omega}^{\text{\,\rm cl}}_{2}.$ (2.5)
Fix $\lambda\in{\Omega}^{\text{\,\rm cl}}_{2}$ and let $\rho$ be as in Lemma 2.3.
If $\operatorname{dist}(f(\lambda),f(K))\geq\rho$ then
$\operatorname{dist}(\lambda,\widetilde{K})\leq\operatorname{diam}\Omega_{2br}\leq(\rho^{-1}\operatorname{diam}\Omega_{2br})\operatorname{dist}(f(\lambda),f(K)).$
Now suppose that $\operatorname{dist}(f(\lambda),f(K))<\rho$. Take a point $w\in f(K)\cap D(f(\lambda),\rho)$ such that $\operatorname{dist}(f(\lambda),f(K))=\lvert f(\lambda)-w\rvert$. Let $g\colon D(f(\lambda),\rho)\to Y_{\lambda}$ be an analytic branch of $f^{-1}$. Since $Y_{\lambda}\subset\Omega_{1}$, we have $g(w)\in\widetilde{K}$. Observe that we also have $\lvert g^{\prime}\rvert\leq\tau^{-1}$ in $D(f(\lambda),\rho)$, where $\tau$ is given by (2.3). Therefore
$\operatorname{dist}(\lambda,\widetilde{K})\leq\lvert g(f(\lambda))-g(w)\rvert\leq\tau^{-1}\lvert f(\lambda)-w\rvert=\tau^{-1}\operatorname{dist}(f(\lambda),f(K)).$
This completes the proof. ∎
For any $z\in\mathbb{C}$, we set $k(z)$ to be the number the number of preimages under $f$ of $z$ in $\Omega_{1br}$, counted with their multiplicities. Notice that there is a uniform bound
$k(z)\leq n_{0},$
because the number of preimages of $z$ on each of the sets $U_{\lambda_{j}},W_{k}$ is uniformly bounded. Let us denote the preimages of $z$ in $\Omega_{1br}$ by $\lambda_{z,1},\dotsc,\lambda_{z,k(z)}$.
For $z\in f(\Omega_{2br})$, define $\varphi_{z}\colon\Omega_{1br}\to\mathbb{C}$ by
$f(\lambda)-z=\varphi_{z}(\lambda)(\lambda-\lambda_{z,1})\dotsb(\lambda-\lambda_{z,k(z)}).$ (2.6)
This function is analytic in $\Omega_{1br}$ and does not vanish there, hence $1/\varphi_{z}$ is also analytic in $\Omega_{1br}$. In the next lemma, we are interested in its restriction to the domain $\Omega_{2br}$, which is smaller than $\Omega_{1br}$.
Lemma 2.5.
We have
$\lVert 1/\varphi_{z}\rVert_{H^{\infty}(\Omega_{2br})}\lesssim 1,\quad z\in f(\Omega_{2br}).$
Proof.
Fix $z\in f(\Omega_{2br})$. Let $\rho$ be the number from Lemma 2.3.
Let $\lambda$ be any point in $\Omega_{2}$ be such that $f(\lambda)\neq z$. If $\lvert f(\lambda)-z\rvert\geq\rho$, then we have the bound
$\left\lvert\frac{1}{\varphi_{z}(\lambda)}\right\rvert=\left\lvert\frac{(\lambda-\lambda_{z,1})\dotsb(\lambda-\lambda_{z,k(z)})}{f(\lambda)-z}\right\rvert\leq\frac{1+(\operatorname{diam}\Omega_{1})^{n_{0}}}{\rho}\,.$
If $\lvert f(\lambda)-z\rvert<\rho$, then there is a neighborhood $Y_{\lambda}\subset\Omega_{1}$ of $\lambda$ such that $Y_{\lambda}\leftrightarrow D(f(\lambda),\rho)$. In particular, for some $j$, the point $\lambda_{z,j}$ belongs to $Y_{\lambda}$. Let $g$ be the analytic branch of $f^{-1}$ from $D(f(\lambda),\rho)$ to $Y_{\lambda}$. We have
$\lvert\lambda-\lambda_{z,j}\rvert$ $=\lvert g(f(\lambda))-g(z)\rvert$
$\leq\Big{[}\sup_{t\in D(f(\lambda),\rho)}\lvert g^{\prime}(t)\rvert\Big{]}\lvert f(\lambda)-z\rvert$
$\leq\tau^{-1}\lvert f(\lambda)-z\rvert$.
Therefore, by (2.6), in this case we obtain the bound
$\left\lvert\frac{1}{\varphi_{z}(\lambda)}\right\rvert\leq\tau^{-1}\big{(}1+(\operatorname{diam}\Omega_{1})^{n_{0}-1}\big{)}.$
The above estimates show that the norms $\|1/\varphi_{z}\|_{H^{\infty}(\Omega_{2})}$ are uniformly bounded for $z\in f(\Omega_{2})$. Since $\partial\Omega_{\text{br}}\subset\Omega_{2}$ and $1/\varphi_{z}$ is analytic on a neighbourhood of ${\Omega}^{\text{\,\rm cl}}_{br}$, the maximum principle gives that $\|1/\varphi_{z}\|_{H^{\infty}(\Omega_{2})}=\|1/\varphi_{z}\|_{H^{\infty}(\Omega_{2br})}$. This completes the proof. ∎
3. Proofs of Theorem A and Theorem 1.1
Proof of Bakaev’s Theorem A.
It will suffice to estimate $\|(f(T)-z)^{-1}\|$ for $z\in f(\Omega_{2br})\setminus f(K)$. Since any such $z$ does not belong to the set $\{f(\xi_{1}),\dotsc,f(\xi_{\ell})\}$, its preimages (under $f$) in $\Omega_{1br}$ are all different. As before, we denote them by $\lambda_{z,1},\dots,\lambda_{z,k(z)}$. Consider the decomposition into simple fractions
$\frac{1}{(\lambda-\lambda_{z,1})\dotsb(\lambda-\lambda_{z,k(z)})}=\frac{c_{z,1}}{\lambda-\lambda_{z,1}}+\dotsb+\frac{c_{z,k(z)}}{\lambda-\lambda_{z,k(z)}}.$
By (2.6),
$(f(\lambda)-z)^{-1}=\varphi_{z}(\lambda)^{-1}\sum_{j=1}^{k(z)}c_{z,j}(\lambda-\lambda_{z,j})^{-1},$
so applying the Riesz–Dunford calculus one gets
$(f(T)-z)^{-1}=\varphi_{z}(T)^{-1}\sum_{j=1}^{k(z)}c_{z,j}(T-\lambda_{z,j})^{-1}.$ (3.1)
By Lemma 2.5, the norms of $\varphi_{z}(T)^{-1}$ are uniformly bounded. Since $k(z)\leq n_{0}$ uniformly in $z$, it will be sufficient to check the estimate for each term in the sum:
$\|c_{z,j}(T-\lambda_{z,j})^{-1}\|\lesssim\operatorname{dist}(z,f(K))^{-s},\quad z\in f(\Omega_{2br})\setminus f(K).$ (3.2)
So, fix an index $j$. Notice that
$|c_{z,j}|=\prod_{k\neq j}|\lambda_{z,j}-\lambda_{z,k}|^{-1}.$ (3.3)
We distinguish two cases.
Case 1: $\lambda_{z,j}\notin\Omega_{\text{br}}$. Then Lemma 2.2 implies that $|c_{z,j}|\leq\max\{1,\delta^{-n_{0}+1}\}$. Therefore, invoking Lemma 2.4 we obtain
$\|c_{z,j}(T-\lambda_{z,j})^{-1}\|$ $\lesssim\operatorname{dist}(\lambda_{z,j},K)^{-s}$
$\lesssim\operatorname{dist}(\lambda_{z,j},\widetilde{K})^{-s}$
$\lesssim\operatorname{dist}(z,f(K))^{-s},\quad z\in f(\Omega_{2br})\setminus f(K).$
Case 2: $\lambda_{z,j}\in\Omega_{\text{br}}$. Hence $\lambda_{z,j}\in W_{k}$ for some $k$. We recall that $W_{k}$ contains a branching point $\xi_{k}$ of $f$ of order $m=m_{k}$. Here we use a part of Bakaev’s arguments. If $\lambda_{z,p}\notin W_{k}$, then $\lvert\lambda_{z,j}-\lambda_{z,p}\rvert\geq\delta$. On the other hand, if $\lambda_{z,p}\in W_{k}$ with $p\neq j$ (there are exactly m1m-1 of these $\lambda_{z,p}$), then
$\lvert\lambda_{z,j}-\lambda_{z,p}\rvert\asymp\lvert\lambda_{z,j}-\xi_{k}\rvert,\quad z\in f(\Omega_{2br})\setminus f(K)$
Therefore
$|c_{z,j}|\asymp|\lambda_{z,j}-\xi_{k}|^{-m+1},\quad z\in f(\Omega_{2br})\setminus f(K).$
Next, let $\zeta_{z,j}$ be a point of $K$ such that $\operatorname{dist}(\lambda_{z,j},K)=\lvert\lambda_{z,j}-\zeta_{z,j}\rvert$. Since $\xi_{k}\in K$, we have $|\zeta_{z,j}-\xi_{k}|\leq 2|\lambda_{z,j}-\xi_{k}|\leq 2\operatorname{diam}W_{j}$. Recall that, by  (2.1), ${D}^{\text{\,\rm cl}}(\xi_{k},2\operatorname{diam}W_{k})\subset\Omega$. Thus we can estimate
$\operatorname{dist}(z,f(K))$ $\leq|f(\lambda_{z,j})-f(\zeta_{z,j})|$
$\leq|\lambda_{z,j}-\zeta_{z,j}|\cdot\max_{{D}^{\text{\,\rm cl}}(\xi_{k},2\operatorname{diam}W_{k})}\,|f^{\prime}|$
$\lesssim\operatorname{dist}(\lambda_{z,j},K)\cdot|\lambda_{z,j}-\xi_{k}|^{m-1},\quad z\in f(\Omega_{2br})\setminus f(K).$
Hence, using that $s\geq 1$ and $m\geq 2$, we have
$\|c_{z,j}(T-\lambda_{z,j})^{-1}\|$ $\lesssim|\lambda_{z,j}-\xi_{k}|^{-m+1}\operatorname{dist}(\lambda_{z,j},K)^{-s}$
$\lesssim|\lambda_{z,j}-\xi_{k}|^{-m+1}\operatorname{dist}(z,f(K))^{-s}|\lambda_{z,j}-\xi_{k}|^{s(m-1)}$
$\lesssim\operatorname{dist}(z,f(K))^{-s},\quad z\in f(\Omega_{2br})\setminus f(K).$
Therefore (3.2) holds, as we wanted to prove. ∎
Lemma 3.1.
If $z\in f(\Omega_{2})\setminus\sigma(f(T))$ and $z\notin f(\Omega_{\text{br}})$, then
$\lVert(f(T)-z)^{-1}\rVert\asymp\max\{\lVert(T-\lambda)^{-1}\rVert:\lambda\in f^{-1}(z)\cap\Omega_{1}\}.$
Proof.
The inequality
$\lVert(f(T)-z)^{-1}\rVert\lesssim\max\{\lVert(T-\lambda)^{-1}\rVert:\lambda\in f^{-1}(z)\cap\Omega_{1}\}$
follows from identities (3.1) and  (3.3). Indeed, since $z\notin f(\Omega_{\text{br}})$, for each $\lambda=\lambda_{z,j}$ satisfying $\lambda\in f^{-1}(\{z\})\cap\Omega_{1}$, Case 1 in the previous proof will take place. Therefore in (3.1), $c_{z,j}$ are uniformly bounded.
To prove the reverse inequality, take any $\lambda\in\Omega_{1}$ and consider the function $\psi_{\lambda}\colon\Omega_{2}\to\mathbb{C}$, given by $\psi_{\lambda}(t)=(f(t)-f(\lambda))/(t-\lambda)$. It is easy to see that the functions $\psi_{\lambda}$ are analytic in $\Omega_{2}$ and their norms in $H^{\infty}(\Omega_{2})$ are uniformly bounded by a finite constant. By the same argument as in Lemma 2.5, the norms of $\psi_{\lambda}$ in $H^{\infty}(\Omega_{2br})$ are uniformly bounded by the same constant. Therefore, given $z\in f(\Omega_{2})\setminus\sigma(f(T))$ and $\lambda\in f^{-1}(z)\cap\Omega_{1}$, the identity
$(T-\lambda)^{-1}=\psi_{\lambda}(T)\big{(}f(T)-z\big{)}^{-1},$
readily gives
$\lVert(T-\lambda)^{-1}\rVert\leq C\lVert(f(T)-z)^{-1}\rVert,$
as we wanted to prove. ∎
Proof of Lemma 1.2.
It is well-known that the function
$u(\lambda):=\log\|(T-\lambda)^{-1}\|$
[…]
Under the hypothesis of (i), AA is contained in a finite union of Lipschitz curves, so that inequality (1.3) and  [3, part (1) of Corollary 2.8] imply that $T$ has resolvent growth of order $s$ near $B=K$.
[…]
Proof of Theorem 1.1.
First we observe that since $f^{\prime}$ does not vanish on $K$, $\Omega_{\text{br}}=\emptyset$ and $\Omega_{2br}=\Omega_{2}$. Suppose $f(T)$ has resolvent growth of order $s$ near $K$. Then, by Lemmas 3.1 and  2.4,
$\lVert(T-\lambda)^{-1}\rVert$ $\lesssim\lVert(f(T)-f(\lambda))^{-1}\rVert$
$\lesssim\operatorname{dist}(f(\lambda),f(K))^{-s}$
$\lesssim\operatorname{dist}(\lambda,\widetilde{K})^{-s},\quad\lambda\in\Omega_{2}.$
First assume that (i) holds, that is, $K$ is contained in a finite union of Lipschitz curves. Then it is easy to see that $\widetilde{K}$ also is contained in a finite union of Lipschitz curves. By applying part (i) of Lemma 1.2, we get that
$\lVert(T-\lambda)^{-1}\rVert\lesssim\operatorname{dist}(\lambda,K)^{-s},\quad\lambda\in\Omega_{2}.$
Now assume that (ii) holds, that is, $K$ is $p$-admissible for some $p\in(0,2)$. Then it is easy to see that $f(K)$ and hence $\widetilde{K}$ also are $p$-admissible. By applying part (ii) of Lemma 1.2, now we get that for any $\sigma>s$,
$\lVert(T-\lambda)^{-1}\rVert\lesssim\operatorname{dist}(\lambda,K)^{-\sigma},\quad\lambda\in\Omega_{2}.\qed$
Example 3.2.
Let $N$ be a compact normal operator on a Hilbert space, whose spectrum is contained in ${D(0,1)}^{\text{\,\rm cl}}$ and is so dense that
$\operatorname{dist}(\lambda,\sigma(N))\leq|\lambda|^{3},\quad|\lambda|\leq 1.$
Let $T=N\oplus J$, where $J$ is the $2\times 2$ Jordan cell corresponding to eigenvalue $3$. Put $f(z)=z^{2}-3z$. Then $f^{\prime}\neq 0$ on $\sigma(T)$ and $T$ fails to have first order resolvent growth near $\sigma(J)=\{3\}$. However, $f(T)$ has first order resolvent growth. Indeed,
$\|(f(T)-zI)^{-1}\|$ $=\max\big{(}\operatorname{dist}(f(N),z)^{-1},\|(f(J)-z)\|^{-1}\big{)}$
$\lesssim\operatorname{dist}(f(N),z)^{-1}$
$\leq\operatorname{dist}(f(T),z)^{-1},\quad|z|\leq 5.$
Indeed, it is easy to see that
$\operatorname{dist}(z,\sigma(N))\lesssim|z|^{3},\quad|z|\leq 1$
and that $f(J)$ is a $2\times 2$ Jordan cell corresponding to eigenvalue 0, so has second order resolvent growth.
This example shows that the geometric requirements on $\sigma(T)$ in Theorem 1.1 cannot be dropped.